# Linear Diffusion Networks: Harnessing Diffusion Processes for Global Interactions

### v2

@@ 4 Theory
In this section, we rigorously establish that the diffusion kernel in LDN inherently captures global dependencies across the entire input sequence. We show that, under mild conditions on the learnable kernel $K$, repeated application of the diffusion update leads to a mixing of information such that every output token is influenced by every input token.
4.1 Global Dependency via Iterated Diffusion
@@
To simplify the discussion, consider the pure diffusion update (ignoring for now the local update $F$ and the linear attention module $A_{\text{lin}}$). At each layer $\ell$, the diffusion module updates the hidden state as
@@
we can write the update as
$H^{(\ell)}=\left(I+\delta t\left(K-\operatorname{diag}(K\mathbf{1})\right)\right)H^{(\ell-1)}.$
Define the effective diffusion operator as
$A\triangleq I+\delta t\Bigl{(}K-\operatorname{diag}(K\mathbf{1})\Bigr{)}.$
After $L$ layers, the cumulative effect of the diffusion process is
@@
4.2 The Global Dependency Theorem
@@
    The directed graph $G(K)$ induced by the nonzero entries of $K$ is strongly connected; that is, for any pair of time indices $i$ and $j$, there exists a sequence $\{i=i_{0},i_{1},\dots,i_{k}=j\}$ with $K_{i_{m+1}i_{m}}>0$ for all $m$. (This is an explicit assumption on the design of $K$; note that while $K$ may contain zero entries, its nonzero pattern must ensure strong connectivity.)
Then, there exists a positive integer $L$ (whose value depends on the structure of $K$) such that the $(i,j)$-th entry of the effective diffusion operator $A^{L}$ is strictly positive for all $i,j$:
$[A^{L}]_{ij}>0.$
Equivalently, the hidden state $h_{i}^{(L)}$ depends on every input $h_{j}^{(0)}$ via
@@
Thus, the diffusion process captures global dependencies.
@@
Starting from the diffusion update,
@@
where
$A\triangleq I+\delta t\left(K-\operatorname{diag}(K\mathbf{1})\right),$
we note that $A$ is a perturbation of the identity matrix by a non-negative matrix. Under the stated assumptions, the induced directed graph is strongly connected, so $A$ is irreducible. By a classical result from the Perron–Frobenius theory for irreducible non-negative matrices, there exists a positive integer $L$ such that $A^{L}$ has strictly positive entries, i.e., $[A^{L}]_{ij}>0$ for all $i,j$.
This shows that every output token $h_{i}^{(L)}$ is influenced by every input token $h_{j}^{(0)}$, as required. ∎
4.3 Implications and Discussion
The theorem formalizes the intuition that even if the diffusion kernel $K$ is designed with local attenuation (for example, via a Gaussian decay $g(|t-s|)$ or directional bias $\phi(t-s)$), the strong connectivity assumption guarantees that there is a nonzero path—possibly through intermediate tokens—between any two tokens in the sequence. However, two important points warrant discussion:
@@ 4.3 Implications and Discussion
    Quantitative Bound on $L$: The theorem guarantees the existence of a positive integer $L$ such that $[A^{L}]_{ij}>0$ for all $i,j$, but it does not provide a practical upper bound on $L$. In applications, if the graph induced by $K$ is only sparsely connected, the required $L$ might be very large (potentially on the order of thousands), which could limit the practical utility of the pure diffusion process in isolation.
@@
    Ensuring Strong Connectivity: Although $K$ is allowed to contain zero entries (as per assumption 1), the theorem’s guarantee relies on the assumption that the pattern of nonzero entries in $K$ induces a strongly connected graph. This is a design or learning constraint: one must ensure that despite local attenuation, there are sufficient nonzero connections such that every token is (indirectly) connected to every other token.
In practice, LDN mitigates these concerns by integrating the diffusion module with both a local update $F$ and a linear attention module $A_{\text{lin}}$. These additional components provide alternative pathways for global information propagation, thereby compensating for scenarios where the required $L$ for the diffusion process alone might be impractically high.
Thus, while the theorem establishes a foundational property of the iterative diffusion process, its practical efficacy in capturing global dependencies within LDN is further enhanced by the complementary design of the overall architecture.
4.4 Stability and Convergence of the Diffusion Process
We now analyze the stability of the pure diffusion update, isolating it from the local and attentional components. Recall that the pure diffusion update is given by
$H^{(\ell)}=\left(I+\delta t\left(K-\operatorname{diag}(K\mathbf{1})\right)\right)H^{(\ell-1)},$
where $K\in\mathbb{R}^{T\times T}$ is the learnable diffusion kernel with $K_{ts}\geq 0$ for all $t,s$ and $\delta t$ is the adaptive time-step parameter.
Theorem (Stability of the Diffusion Update):
Assume that the kernel $K$ satisfies $K_{ts}\geq 0$ for all $t,s$ and that the time-step $\delta t$ is chosen such that
@@ 4.4 Stability and Convergence of the Diffusion Process
Then, for each layer $\ell$, the diffusion update is a convex combination of the hidden states at layer $\ell-1$. In particular, for every $t$,
@@
with coefficients satisfying
@@
Consequently, the diffusion update is non-expansive with respect to the $\ell_{2}$ norm:
@@
Repeated application of this update thus ensures that the propagation of hidden states remains stable and does not amplify perturbations.
@@
For each time step $t$, rewrite the update as
@@
$\alpha_{t}\triangleq 1-\delta t\sum_{s=1}^{T}K_{ts}.$
By the choice of $\delta t$, we have $\alpha_{t}\geq 0$ and the sum of the coefficients is
$\alpha_{t}+\delta t\sum_{s=1}^{T}K_{ts}=1.$
Thus, $h_{t}^{(\ell)}$ is a convex combination of $\{h_{s}^{(\ell-1)}\}_{s=1}^{T}$. Because convex combinations are non-expansive with respect to the $\ell_{2}$ norm, it follows that
$\|h_{t}^{(\ell)}\|_{2}\leq\max_{s}\|h_{s}^{(\ell-1)}\|_{2}.$
Aggregating over all time steps yields
$\|H^{(\ell)}\|_{2}\leq\|H^{(\ell-1)}\|_{2}.$
This completes the proof. $\blacksquare$
@@
This stability guarantee is crucial for the practical training of LDN. By ensuring that the hidden state norms do not increase with layer depth, the theorem alleviates concerns over exploding gradients and uncontrolled error propagation. In tandem with the global mixing properties established earlier, this result underpins the balanced design of LDN—merging robust stability with comprehensive global information propagation.
5 Experiments
In this section, we evaluate LDN on multiple benchmarks, comparing its performance against the Vision Transformer (ViT) baseline Dosovitskiy et al. (2020). We report results on both the large-scale ImageNet dataset Deng et al. (2009) and the CIFAR-10 dataset cif . In addition, we assess the performance of LDN on the GLUE benchmark Wang et al. (2018) and conduct ablation studies to examine the contributions of LDN’s individual components.
5.1 Datasets and Experimental Setup
For ImageNet, we follow standard preprocessing and augmentation strategies as described in Deng et al. (2009). For CIFAR-10, we adopt typical data augmentation techniques (e.g., random cropping and horizontal flipping) to improve generalization. All models are trained using the Adam optimizer Kingma and Ba (2014) with a cosine annealing learning rate schedule, and experiments were conducted on NVIDIA GPUs under comparable hardware settings.
Pretraining for GLUE.
In order to achieve competitive performance on natural language tasks, LDN is first pretrained on a large-scale unlabeled corpus. Following the setups of BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019), we combine English Wikipedia and the BookCorpus to form our pretraining dataset. We employ a masked language modeling task augmented with a novel diffusion-based objective that iteratively refines token representations by propagating contextual information. This approach is designed to capture both local and global dependencies effectively. Pretraining is carried out for 1 million steps with a batch size of 256, resulting in a model of approximately 125 million parameters—comparable to both BERT-base and RoBERTa-base—to ensure a fair comparison during fine-tuning on downstream GLUE tasks.
5.2 Results on ImageNet
Table 1 summarizes the performance comparison between ViT and LDN across different model scales (Base, Large, and Huge). Notice that LDN not only achieves higher Top-1 and Top-5 accuracies but does so with fewer parameters and lower computational cost (FLOPs).
Table 1: Comparison of ViT and LDN on ImageNet.
Variant Architecture Params (M) FLOPs (GMac) Top-1 (%) Top-5 (%)
Base ViT 86 17.6 82.5 96.0
LDN 52 10.6 83.1 96.6
Large ViT 307 61.6 84.2 97.0
LDN 181 36.7 84.8 97.3
Huge ViT 632 120.5 84.8 97.3
LDN 373 75.4 85.3 97.8
5.3 Results on CIFAR-10
We further evaluate the models on CIFAR-10. As shown in Table 2, LDN achieves a test accuracy of 80.9%, outperforming the ViT baseline which attains 77.5%. Figure 2 plots the training accuracy over epochs, illustrating that LDN converges faster and reaches a higher accuracy than ViT.
Table 2: Test Accuracy on CIFAR-10.
Model Test Accuracy (%)
ViT 77.5
LDN 80.9
Figure 2: Training accuracy comparison between ViT and LDN on CIFAR-10.
5.4 Results on GLUE Benchmark
To assess LDN’s versatility in natural language processing, we fine-tune the pretrained model on several tasks from the GLUE benchmark Wang et al. (2018). Table 3 compares LDN against popular methods such as BERT-base Devlin et al. (2019) and RoBERTa-base Liu et al. (2019). LDN demonstrates competitive performance across multiple GLUE tasks, particularly excelling on tasks that require complex reasoning.
Table 3: Performance on GLUE benchmark tasks.
Model CoLA SST-2 MRPC MNLI QNLI Avg.
BERT-base 52.1 93.2 88.9 84.6 90.5 81.8
RoBERTa-base 60.1 94.5 90.3 86.0 91.7 84.5
LDN 62.3 95.1 91.2 87.4 92.3 85.6
5.5 Ablation Studies
To understand the individual contributions of LDN’s components (diffusion module, local update, and linear attention), we perform an ablation study on CIFAR-10. Table 4 reports the test accuracy when selectively removing one or more components. The full model (with all three components) achieves the highest accuracy, confirming that each module is critical to the overall performance.
Table 4: Ablation study on CIFAR-10 components.
Model Variant Diffusion Local Update Linear Attention Test Acc (%)
Full LDN 80.9
Without Diffusion 78.0
Without Local Update 79.2
Without Linear Attention 79.7
ViT Baseline 77.5
5.6 Discussion
The experimental results demonstrate that LDN consistently outperforms the ViT baseline on both ImageNet and CIFAR-10. On ImageNet, LDN not only achieves higher Top-1 and Top-5 accuracies but does so with reduced model complexity and computational cost. On CIFAR-10, the enhanced test accuracy and faster convergence—as illustrated by the training curves in Figure 2—underscore the effectiveness of the integrated diffusion process in capturing global dependencies.
Furthermore, the evaluation on the GLUE benchmark (Table 3) indicates that LDN is competitive with, and in some cases exceeds, the performance of popular NLP models such as BERT-base and RoBERTa-base. Our ablation study (Table 4) confirms that each module within LDN contributes significantly to the overall performance, highlighting the importance of a unified architecture that leverages both local and global information.
Overall, these experiments confirm the benefits of reinterpreting temporal (or spatial) information sharing as a diffusion process. The promising results open up avenues for further research, such as extending the model to larger datasets and exploring enhanced non-linear attention mechanisms. Future work will investigate adaptive strategies for optimizing the diffusion kernel parameters and incorporating more sophisticated integration schemes to further improve both efficiency and performance.
6 Conclusion
Empirical evaluations on ImageNet, CIFAR-10, and the GLUE benchmark demonstrate that LDN outperforms competitive baselines such as ViT Dosovitskiy et al. (2020), BERT Devlin et al. (2019), and RoBERTa Liu et al. (2019). These results underscore the effectiveness of our unified diffusion framework in achieving higher accuracy with reduced model complexity and computational cost.

### v4

@@ Appendix A Theory
In this section, we demonstrate rigorously that the diffusion kernel in LDN captures global dependencies across the input sequence. Under mild conditions on the learnable kernel $K$, repeated application of the diffusion update leads to an effective mixing, ensuring that every output token is influenced by every input token.
A.1 Global Dependency via Iterated Diffusion
@@
For clarity, consider the pure diffusion update, setting aside the local update $F$ and the linear attention module $A_{\text{lin}}$. At each layer $\ell$, the hidden state is updated as:
@@
this becomes:
$H^{(\ell)}=\Bigl{(}I+\delta t\Bigl{(}K-\operatorname{diag}(K\mathbf{1})\Bigr{)}\Bigr{)}H^{(\ell-1)}.$
Defining the effective diffusion operator as:
$A\triangleq I+\delta t\Bigl{(}K-\operatorname{diag}(K\mathbf{1})\Bigr{)},$
after $L$ layers the process yields:
@@
A.2 The Global Dependency Theorem
@@
    The directed graph $G(K)$ induced by the nonzero entries of $K$ is strongly connected; that is, for any two time indices $i$ and $j$, there exists a sequence $\{i=i_{0},i_{1},\dots,i_{k}=j\}$ with $K_{i_{m+1}i_{m}}>0$ for all $m$.
Then, there exists a positive integer $L$ (dependent on the structure of $K$) such that every entry of the effective diffusion operator satisfies:
$[A^{L}]_{ij}>0\quad\text{for all }i,j.$
Equivalently, every output token $h_{i}^{(L)}$ depends on every input token $h_{j}^{(0)}$:
@@
Thus, the diffusion process inherently captures global dependencies.
@@
Starting from
@@
with
$A=I+\delta t\Bigl{(}K-\operatorname{diag}(K\mathbf{1})\Bigr{)},$
observe that $A$ is a perturbation of the identity matrix by a non-negative matrix. Given the strong connectivity of the directed graph corresponding to $K$, $A$ is irreducible. By the Perron–Frobenius theorem for irreducible non-negative matrices, there exists a positive integer $L$ such that all entries of $A^{L}$ are strictly positive. This completes the proof. ∎
A.3 Discussion
This theorem formalizes the intuition that even if the diffusion kernel $K$ employs local attenuation (for instance, through a Gaussian decay $g(|t-s|)$ or a directional bias $\phi(t-s)$), the strong connectivity ensures that every token is indirectly linked to every other token. Two points merit further consideration:
@@ A.3 Discussion
    Quantitative Bound on $L$: While the theorem guarantees an $L$ exists such that $[A^{L}]_{ij}>0$ for all $i,j$, it does not provide a practical bound. In scenarios where the graph induced by $K$ is sparse, $L$ might be large, which may affect the efficiency of the diffusion process.
@@
    Ensuring Strong Connectivity: The design of $K$ must ensure that its nonzero pattern yields a strongly connected graph. This is a crucial requirement for achieving global dependency.
In practice, LDN compensates for potential limitations of the pure diffusion process by combining it with a local update $F$ and a linear attention module $A_{\text{lin}}$, thereby offering multiple pathways for global information flow.
A.4 Stability and Convergence of the Diffusion Process
Next, we examine the stability of the pure diffusion update in isolation. Recall that:
$H^{(\ell)}=\left(I+\delta t\Bigl{(}K-\operatorname{diag}(K\mathbf{1})\Bigr{)}\right)H^{(\ell-1)},$
where $K\in\mathbb{R}^{T\times T}$ is a learnable kernel with $K_{ts}\geq 0$ for all $t,s$, and $\delta t$ is the adaptive time-step.
Stability Theorem:
Assume $K_{ts}\geq 0$ for all $t,s$ and choose $\delta t$ such that:
@@ A.4 Stability and Convergence of the Diffusion Process
Then, for every layer $\ell$ and each time step $t$, the update can be written as:
@@
with the coefficients satisfying:
@@
Thus, the update is a convex combination of the hidden states from the previous layer, implying:
@@
Expressing the update for each $t$ as:
@@
$\alpha_{t}\triangleq 1-\delta t\sum_{s=1}^{T}K_{ts},$
and noting that $\alpha_{t}\geq 0$ with $\alpha_{t}+\delta t\sum_{s=1}^{T}K_{ts}=1$, it follows that each $h_{t}^{(\ell)}$ is a convex combination of $\{h_{s}^{(\ell-1)}\}_{s=1}^{T}$. Since convex combinations are non-expansive with respect to the $\ell_{2}$ norm, the result holds. $\blacksquare$
@@
This stability result is vital for training LDN. It ensures that the norms of the hidden states do not increase with each layer, preventing issues like exploding gradients. Together with the global mixing properties, this underlines the balanced design of LDN, combining robustness with effective global information propagation.